OpenStudy (anonymous):

Complete the square. Tutor confused me y=2x^2+6x+2 ->y=-2(x^2-3x)+2 (I'm not sure how or why he got 9/2 in this part) ->y=-2(x^2-3x+9/4)+2+9/2 Final answer: y=-2(x-3/2)^2+13/2 [this is for transformations and professor wants us to put it in standard form before we graph]

OpenStudy (anonymous):

well since he subtracted 9/2 he added it cause he cant really change anything -2 * 9/4 = -9/2 so you must add 9/2

OpenStudy (anonymous):

how did it get from [x^2-3x+9/4] to [x-3/2]^2 not really sure how it was simplified

OpenStudy (anonymous):

whenever you complete square the method is : x^2 - 3x so 3 is 2ab in this case a = 1 b = ? -3 = 2b b = -3/2 now look at what happened : a^2x^2 - abx + b^2 = (ax+b)^2 a =1 b = 3/2 -> (x -3/2)^2

OpenStudy (anonymous):

Thank you =]